# ID-Composer: Multi-Subject Video Synthesis with Hierarchical Identity Preservation

### v2

@@ Abstract
Video generative models pretrained on large-scale datasets can produce high-quality videos, but are often conditioned on text or a single image, limiting controllability and applicability. We introduce ID-Composer, a novel framework that addresses this gap by tackling multi-subject video generation from a text prompt and reference images. This task is challenging as it requires preserving subject identities, integrating semantics across subjects and modalities, and maintaining temporal consistency. To faithfully preserve the subject consistency and textual information in synthesized videos, ID-Composer designs a hierarchical identity-preserving attention mechanism, which effectively aggregates features within and across subjects and modalities. To effectively allow for the semantic following of user intention, we introduce semantic understanding via pretrained vision-language model (VLM), leveraging VLM’s superior semantic understanding to provide fine-grained guidance and capture complex interactions between multiple subjects. Considering that standard diffusion loss often fails in aligning the critical concepts like subject ID, we employ an online reinforcement learning phase to drive the overall training objective of ID-Composer into RLVR. Extensive experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality. Code and training data will be released.
Figure 1: Given a text prompt and multiple reference images, ID-Composer generates subject-consistent videos and achieves impressive subject ID preservation compared with the previous state-of-the-art methods, such as Phantom (Liu et al., 2025b). Kindly zoom in for details.
@@ 1 Introduction
Recent advances in video generative models (OpenAI, 2023; Yang et al., 2024b; Vidu, 2024; keling, 2024; Wan, 2025) achieve high-fidelity results, yet their reliance on sparse inputs, typically a text prompt or a single initial frame, severely restricts controllability. Inspired by the compositional flexibility of modern image generators (Deng et al., 2025a; Fortin et al., 2025), we address this limitation by focusing on multi-subject video generation. This capability further facilitates downstream applications in subject-driven video synthesis, dynamic scene composition, and controllable product placement, and promises to advance personalized content creation, virtual storytelling, and advertising. Despite its impact, multi-subject video generation still faces challenging issues compared to the existing pipeline established for generating single-subject video: (1) Identity preservation is increasingly difficult for multiple subjects throughout the video; (2) Semantic integration on more subjects and modalities involves challengingly aligning and balancing semantic information; (3) Temporal consistency of multi-subject video demands preserving coherent appearance and motion across consecutive frames. To address it, a naive solution is to develop an agentic system that first leverages off-the-shelf subject-driven image generation models (Ruiz et al., 2023) to generate the first frame given the text prompt and reference images, and then applies video generative models (OpenAI, 2023; Yang et al., 2024b; Wan, 2025) to produce the subsequent frames. However, this two-stage pipeline lacks explicit correspondence between the video and the reference subjects, and heavily depends on the quality of the initial frame, resulting in suboptimal identity preservation, poor temporal consistency, and vulnerability to error accumulation. To tackle these challenges, recent works (Deng et al., 2025b; Liu et al., 2025b; Xue et al., 2025a) propose end-to-end solutions by injecting information of multiple subjects into the Multimodal Diffusion Transformers (MMDiT) (Esser et al., 2024) of pretrained video diffusions (Wan, 2025). Although effective, these methods simply concatenate visual tokens of all subjects and textual tokens to noise tokens. This results in higher perplexity during semantic integration and conflict resolution across different subjects, as well as suboptimal performance due to lacking explicit modeling of interactions among subjects and modalities.
To this end, we propose ID-Composer, a novel framework for multi-subject video generation that achieves state-of-the-art performance in identity preservation, temporal consistency, and overall video quality. The key designs of ID-Composer are twofold: (1) A hierarchical identity-preserving attention mechanism. The task we are addressing involves multiple subjects and modalities, making it crucial to effectively aggregate features both within and across different conditions. Inspired by recent work using local-global attention (Wang et al., 2025a; Lin et al., 2025), we design a hierarchical attention mechanism that performs attention on three progressive levels: intra-subject, inter-subject, and cross-modal attention. This design enables effective identity consistency and textual faithfulness by capturing fine-grained details within each subject, modeling interactions among multiple subjects, and integrating information across modalities. (2) Semantic understanding via pretrained vision-language model (VLM). Our model leverages VLM as a text encoder to provide fine-grained guidance and capture complex interactions between multiple subjects. Compared to traditional text encoders (Radford et al., 2021), VLMs (Yang et al., 2024a) demonstrate superior semantic understanding and cross-modal alignment capabilities, which are crucial for the multi-subject video generation. ID-Composer is built upon a strong video diffusion model (Wan, 2025) and incorporates the above two designs into the MMDiT (Esser et al., 2024) architecture, enabling effective generation of high-quality videos with multiple specific subjects. To incentivize the model to generate videos with high quality and strong identity preservation, we further introduce an online reinforcement learning (Liu et al., 2025a; Guo et al., 2025) phase subsequent to the initial diffusion training. A new dataset is constructed to facilitate the training and evaluation of multi-subject video generation models. We introduce a data-curation pipeline leveraging vision-language models and advanced image editing models. Our process involves automated filtering, captioning, and subject-decomposing stages to process video clips, resulting in a mixed-modality dataset.
As shown in Fig. 1, ID-Composer is capable of generating high-quality videos that accurately preserve the identities of multiple subjects, align with the prompt, and maintain temporal consistency. Experiments demonstrate that our model surpasses existing methods in identity preservation, temporal consistency, and video quality. Ablation studies validate the effectiveness of our proposed designs.
In summary, our contributions are as follows:
  • We propose ID-Composer, a novel framework for multi-subject video generation that effectively preserves identities, integrates semantics, and maintains temporal consistency.
  • Our proposed hierarchical identity-preserving attention mechanism, the use of pretrained VLM for semantic understanding, and the incorporation of online reinforcement learning significantly enhance the model’s performance in handling multiple subjects and modalities.
  • We construct a new dataset for multi-subject video generation, and conduct extensive experiments demonstrating the superiority of ID-Composer over existing methods.
@@ 2 Related Works
Recent advances in subject-driven image generation have enabled the creation of high-fidelity visual assets, profoundly impacting digital content creation. A prominent line of work involves tuning-based methods, which achieve high subject fidelity by fine-tuning a generative model on a small set of reference images. These approaches range from optimizing the entire model’s weights, as in DreamBooth (Ruiz et al., 2023), to learning subject-specific textual embeddings (Gal et al., 2022) or adopting parameter-efficient fine-tuning strategies (Hu et al., 2021; Han et al., 2023; Yuan et al., 2023). However, a significant drawback of these methods is the substantial computational overhead required for per-subject optimization. To circumvent this limitation, tuning-free methods (Ye et al., 2023; Wang et al., 2024; Sun et al., 2024; Wu et al., 2024; Liu et al., 2025c) have been proposed, which inject identity information during inference via various conditioning mechanisms, thereby offering a more efficient alternative. More recently, a new paradigm has emerged with versatile, unified models (Chen et al., 2024; Xiao et al., 2024; Lu, 2024; Wu et al., 2025a; Gao et al., 2025; Fortin et al., 2025) that incorporate subject-driven generation as one of many capabilities. Despite this progress, extending subject-driven generation from the image to the video domain remains a formidable challenge, primarily due to the difficulty of maintaining both identity and temporal consistency across frames.
@@
Subject-consistent video generation (keling, 2024; Vidu, 2024; Pika, 2024; Jiang et al., 2025) is commonly achieved by enhancing the model’s attention mechanisms to incorporate appearance cues. More flexible paradigms have also emerged, including adapter-based methods like ID-Animator (He et al., 2024), ConsisID (Yuan et al., 2024), and Stand-In (Xue et al., 2025a). For the specific challenge of handling multiple subjects, a router mechanism has been introduced (Mou et al., 2025). Other works have explored richer customization dimensions: Wang et al. (2025b) design FantasyPortrait to enhance multi-character portrait animation through expression-augmented diffusion transformers. Hu et al. (2025) propose PolyVivid to improve vividness and consistency in multi-subject scenarios via cross-modal interaction. Huang et al. (2025) present Videomage, which enables both multi-subject and motion customization in text-to-video diffusion. Our work is most closely related to Phantom (Liu et al., 2025b), Concat-ID (Zhong et al., 2025), SkyReels-A2 (Fei et al., 2025), and CINEMA (Deng et al., 2025b), which focus on attention-based feature injection. However, these methods, as well as the above extensions, often struggle with conflicts between subject identity and textual prompts and with maintaining consistency. Moreover, ID-Composer proposes a novel method to mitigate these conflicts and enforce multi-level consistency, thereby improving the robustness and quality of multi-subject video generation.
@@
Leveraging its computational efficiency, Direct Preference Optimization (DPO) (Gao et al., 2025) has become a prevalent technique for post-training text-to-image models to align them with human preferences, such as aesthetics and prompt fidelity. Subsequent work, like DenseDPO (Wu et al., 2025b), extends this paradigm to the video domain by refining paired data construction and the granularity of preference labels, yielding videos with superior visual quality and motion dynamics. A primary limitation of these methods, however, is their offline nature, which precludes online parameter updates. This challenge has been recently addressed by verifiable reward-based models like OpenAI’s o1 (Jaech et al., 2024) and Deepseek-R1 (Guo et al., 2025). Concurrently, the growing interest in online learning has spurred various online Reinforcement Learning (RL) methods, including ReFL (Xu et al., 2023), DRaFT (Clark et al., 2023), FlowGRPO (Liu et al., 2025a), DanceGRPO (Xue et al., 2025b), and SRPO (Shen et al., 2025), which have demonstrated promising results in enhancing image aesthetics and text-rendering fidelity. Nevertheless, their application to subject-consistent video generation remains limited. To bridge this gap, we introduce task-specific rewards tailored for subject-consistent video generation, a method that mitigates reward hacking and delivers videos with high identity consistency.
Figure 2: ID-Composer Overview. Our model incorporates a hierarchical identity-preserving attention mechanism and a VLM that performs reasoning on the multimodal input into a video DiT to enable multi-subject video generation. An online RL stage further refines the concept alignment.
@@ 3 Method
3.1 Preliminaries: Rectified Flow for Video Generation
Our methodology is founded upon a latent video diffusion transformer (Peebles & Xie, 2023; Wan, 2025) trained with Rectified Flow (RF) (Lipman et al., 2022; Liu et al., 2022; Esser et al., 2024). In latent video diffusion, a video clip $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is first projected into a latent space by the encoder $\mathcal{E}$ of a variational autoencoder (VAE), yielding the representation $\mathbf{z}_{0}:=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{f\times c\times h\times w}$. The corresponding VAE decoder is denoted by $\mathcal{D}(\mathbf{z}_{0})$. The generative process is formulated based on Rectified Flow, which defines a straight-line trajectory between a sample from the data distribution $\mathbf{z}_{0}\sim p_{\text{data}}$ and a sample from a simple prior distribution, typically a standard Gaussian $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The trajectory is parameterized by a time variable $t\in[0,1]$:
$\mathbf{z}_{t}:=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}.$ (1)
@@ 3.1 Preliminaries: Rectified Flow for Video Generation
$\mathcal{L}_{\text{RF}}:=\mathbb{E}_{t,\mathbf{z}_{0},\bm{\epsilon}}\left[w(t)\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{C}_{\text{ctx}})-(\bm{\epsilon}-\mathbf{z}_{0})\right\rVert_{2}^{2}\right],$ (2)
@@
3.2 ID-Composer: Multi-Subject Video Generation
@@
As shown in Fig. 2, ID-Composer extends the DiT-based latent video diffusion architecture by introducing two key architectural innovations to handle the complexities of multi-subject conditioning: (1) A hierarchical identity-preserving attention mechanism, which is designed to hierarchically aggregate features. (2) Semantic understanding via a pretrained Vision-Language Model (VLM), which we employ as a sophisticated encoder for both textual and visual inputs. To further incentivize the model’s semantic alignment and identity preservation, we incorporate an online reinforcement learning phase subsequent to the initial flow matching training.
3.2.1 Model Architecture
Hierarchical Identity-Preserving Attention.
@@ 3.2.1 Model Architecture
Our model is based on multi-modal DiT (MMDiT), where latent video tokens and conditioning tokens are processed together through a series of transformer blocks. Following this, we first encode the input reference images $\mathcal{I}$ using a pretrained image encoder (e.g., Wan-VAE) to obtain visual feature maps $\{\mathbf{F}_{k}\}_{k=1}^{N}$, where each $\mathbf{F}_{k}\in\mathbb{R}^{c\times h\times w}$ corresponds to the $k$-th subject. These feature maps are then flattened into token sequences $\{\mathbf{f}_{k}\}_{k=1}^{N}$, with each $\mathbf{f}_{k}\in\mathbb{R}^{hw\times c}$. The textual prompt $\mathbf{C}_{\text{txt}}$ is encoded using a pretrained text encoder to produce a sequence of text tokens $\mathbf{f}_{\text{txt}}\in\mathbb{R}^{l\times c}$, where $l$ is the number of text tokens. The complete conditioning token set is thus $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]$. To effectively integrate information from multiple subjects and modalities, we introduce a hierarchical attention mechanism within each transformer block. This mechanism operates in three stages: (1) Intra-subject attention: Each subject’s tokens $\mathbf{f}_{k}$ undergo self-attention independently, allowing the model to capture fine-grained details and spatial relationships within each subject. (2) Inter-subject attention: The outputs from the intra-subject attention are concatenated and subjected to another self-attention layer, enabling the model to learn interactions and relationships between different subjects. (3) Multi-modal attention: Finally, the combined subject features are concatenated with the text tokens $\mathbf{f}_{\text{txt}}$ and the latent video tokens, followed by a self-attention operation that fuses information across modalities. This hierarchical approach ensures that the model can effectively preserve individual subject identities while also capturing complex inter-subject dynamics and aligning them with the textual prompt.
Semantic Reasoning via Vision-Language Models.
While the hierarchical attention mechanism is designed to preserve identity, composing multiple subjects coherently in accordance with ambiguous user instructions remains a formidable challenge. Furthermore, models trained on limited data often lack the reasoning capabilities required for robust generalization to diverse real-world scenarios. To address these limitations, we leverage a powerful pretrained Vision-Language Model (VLM) to achieve a more sophisticated semantic understanding of the prompt and guide the composition of multiple subjects. We employ the Qwen2.5-VL model (Bai et al., 2025) for its superior multimodal reasoning capabilities. The VLM processes both the textual prompt $\mathbf{C}_{\text{txt}}$ and the set of reference images $\mathcal{I}$ to generate semantically enriched text tokens, which we denote as $\mathbf{f}_{\text{txt}}$: $\mathbf{f}_{\text{txt}}=\text{VLM}_{\text{enc}}(\mathbf{C}_{\text{txt}},\mathcal{I})\in\mathbb{R}^{l^{\prime}\times c}.$ These tokens capture not only the textual semantics but also the visual concepts of the subjects, enabling a more holistic understanding of the user’s intent. These VLM-enhanced tokens then replace the embeddings from a standard text encoder. The complete set of conditioning tokens $\mathbf{C}_{\text{ctx}}$ is subsequently formed by concatenating these enhanced text tokens with the individual subject tokens: $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]\in\mathbb{R}^{(l^{\prime}+N\cdot hw)\times c}.$ By integrating the VLM’s advanced reasoning, ID-Composer achieves superior semantic alignment and identity preservation, resulting in higher-quality videos that faithfully adhere to complex, multi-subject prompts.
3.3 Training with Online Reinforcement Learning
GRPO
We use Flow-GRPO (Liu et al., 2025a) to perform online reinforcement learning (RL) to further improve the video generation quality. For each condition $q$, GRPO (Guo et al., 2025) samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective:

### v3

# ID-Crafter: VLM-Grounded Online RL for Compositional Multi-Subject Video Generation

@@ Abstract
Significant progress has been achieved in high-fidelity video synthesis, yet current paradigms often fall short in effectively integrating identity information from multiple subjects. This leads to semantic conflicts and suboptimal performance in preserving identities and interactions, limiting controllability and applicability. To tackle this issue, we introduce ID-Crafter, a framework for multi-subject video generation that achieves superior identity preservation and semantic coherence. ID-Crafter integrates three key components: (i) a hierarchical identity-preserving attention mechanism that progressively aggregates features at intra-subject, inter-subject, and cross-modal levels; (ii) a semantic understanding module powered by a pretrained Vision-Language Model (VLM) to provide fine-grained guidance and capture complex inter-subject relationships; and (iii) an online reinforcement learning phase to further refine the model for critical concepts. Furthermore, we construct a new dataset to facilitate robust training and evaluation. Extensive experiments demonstrate that ID-Crafter establishes new state-of-the-art performance on multi-subject video generation benchmarks, excelling in identity preservation, temporal consistency, and overall video quality.
Figure 1: Given a text prompt and multiple reference images, ID-Crafter generates subject-consistent videos and achieves impressive subject ID preservation (e.g., face score) compared with the previous state-of-the-art methods, such as Phantom [liu2025phantom].
@@ 1 Introduction
Recent advances in video generative models [OpenAI2023Sora, yang2024cogvideox, Vidu, Keling, wan2025] have achieved remarkable fidelity video generation, yet their primary reliance on sparse inputs like a single text prompt or initial frame significantly limits their controllability, especially for complex scenes. Drawing inspiration from the compositional flexibility of modern image generators [deng2025emergingpropertiesunifiedmultimodal, nanobanana], we focus on the challenging task of multi-subject video generation. This capability is crucial for downstream applications such as subject-driven video synthesis, dynamic scene composition, and controllable product placement, promising to advance personalized content creation, virtual storytelling, and advertising.
However, extending existent single-subject video generation techniques [he2024id, yuan2024identity, xue2025standinlightweightplugandplayidentity, sang2025lynx] to multiple subjects introduces a significant challenge: maintaining the distinct identity of each subject while ensuring temporal coherent throughout the video. Recent studies [CINEMA, liu2025phantom] attempt to address this by injecting multi-subject information into pretrained video diffusion models [wan2025]. But these methods struggle to resolve semantic conflicts between subjects and leads to a degradation of identity. The fundamental challenge lies in the inherent tension between preserving individual subject identities and generating a coherent, dynamic scene—a conflict that current methods fail to resolve effectively.
To address this dilemma, we propose ID-Crafter, which harmonizes identity consistency with motion quality by integrating three strategic components. ID-Crafter is specifically engineered to tackle the inherent challenges of multi-subject video synthesis: (1) First, to disentangle complex inter-subject relationships while preserving individual attributes, we introduce a hierarchical identity-preserving attention mechanism. This module operates in a cascaded manner: it first attends to intra-subject details to capture unique features, subsequently models inter-subject interactions to mitigate identity leakage, and finally performs cross-modal attention to ensure semantic alignment with the textual prompt. This hierarchical structure establishes a robust foundation for identity preservation. (2) Second, to enhance generation fidelity for complex descriptions, we utilize a foundational VLM [Qwen2.5-VL] for semantic feature extraction. unlike conventional text encoders [radford2021learning], VLMs offer a granular understanding of scene composition and cross-modal dynamics, which is essential for accurately interpreting and rendering intricate multi-subject scenarios. (3) Third, to directly optimize the trade-off between identity preservation, video quality, and motion fluidity, we introduce an online Reinforcement Learning (RL) strategy. Optimizing a multi-subject diffusion policy presents significant stability challenges. Inspired by Flow-GRPO [flowgrpo], we address this by employing groupwise comparison for advantage estimation [shao2024deepseekmath, Deepseek-R1], thereby eliminating the need for a fragile value network. Since applying Online RL (e.g., ReFL [ReFL] or DRaFT [DRaFT]) to a multi-subject generative architecture presents non-trivial challenges, as the optimization policy encompasses the video diffusion model augmented with our hierarchical ID attention. We address this by utilizing groupwise comparison for advantage estimation [shao2024deepseekmath, Deepseek-R1], thereby eliminating the reliance on a fragile value network. Furthermore, we introduce a contrastive learning mechanism that provides a stable training signal to fine-tune the hierarchical attention, directly maximizing identity fidelity while mitigating reward hacking.
As illustrated in Fig. 1, these synergistic components empower ID-Crafter to generate high-fidelity videos distinguished by precise identity preservation, robust prompt alignment, and exceptional temporal consistency. Consequently, our method markedly surpasses previous methods (e.g., Phantom [liu2025phantom]), in producing subject-consistent videos. Our contributions are summarized as follows:
  • We propose ID-Crafter, a unified framework that mitigates the tension between identity preservation and motion fluidity in multi-subject video generation through the integration of hierarchical attention, VLM-based semantic guidance, and RL-based fine-tuning.
  • We pioneer the application of online RL to video generation, demonstrating its effectiveness and stability in optimizing complex perceptual rewards and balancing competing generation objectives.
  • Extensive experiments show the superiority of proposed method over existing methods in terms of identity preservation, temporal consistency, and overall video quality.
@@ 2 Related Works
Recent advances in subject-driven image generation have enabled the creation of high-fidelity visual assets, profoundly impacting digital content creation. A prominent line of work involves tuning-based methods, which achieve high subject fidelity by fine-tuning a generative model on a small set of reference images. These approaches range from optimizing the entire model’s weights, as in DreamBooth [ruiz2023dreambooth], to learning subject-specific textual embeddings [gal2022image] or adopting parameter-efficient fine-tuning strategies [hu2021lora, han2023svdiff, yuan2023inserting]. However, a significant drawback of these methods is the substantial computational overhead required for per-subject optimization. To circumvent this limitation, tuning-free methods [ye2023ip, wang2024instantid, sun2024rectifid, wu2024infinite, liu2025mosaic] have been proposed, which inject identity information during inference via various conditioning mechanisms, thereby offering a more efficient alternative. More recently, a new paradigm has emerged with versatile, unified models [chen2024unireal, xiao2024omnigen, erwold-2024-qwen2vl-flux, wu2025qwen, seedream, nanobanana] that incorporate subject-driven generation as one of many capabilities. Despite this progress, extending subject-driven generation from the image to the video domain remains a formidable challenge, primarily due to the difficulty of maintaining both identity and temporal consistency across frames.
@@
Subject-consistent video generation [Keling, Vidu, Pika, VACE] is commonly achieved by enhancing the model’s attention mechanisms to incorporate appearance cues. More flexible paradigms have also emerged, including adapter-based methods like ID-Animator [he2024id], CustomVideo [wang2024customvideo], ConsisID [yuan2024identity], and Stand-In [xue2025standinlightweightplugandplayidentity]. For the specific challenge of handling multiple subjects, a router mechanism has been introduced [mou2025dreamo]. Other works have explored richer customization dimensions: wang2025fantasyportrait design FantasyPortrait to enhance multi-character portrait animation through expression-augmented diffusion transformers. hu2025polyvivid propose PolyVivid to improve vividness and consistency in multi-subject scenarios via cross-modal interaction. huang2025videomage present Videomage, which enables both multi-subject and motion customization in text-to-video diffusion. Our work is most closely related to Phantom [liu2025phantom], Concat-ID [Concat-ID], SkyReels-A2 [fei2025skyreels], and CINEMA [CINEMA], which focus on attention-based feature injection. However, these methods, as well as the above extensions, often struggle with conflicts between subject identity and textual prompts and with maintaining consistency. Moreover, ID-Crafter proposes a novel method to mitigate these conflicts and enforce multi-level consistency, thereby improving the robustness and quality of multi-subject video generation.
@@
Leveraging its computational efficiency, Direct Preference Optimization (DPO) [seedream] has become a prevalent technique for post-training text-to-image models to align them with human preferences, such as aesthetics and prompt fidelity. Subsequent work, like DenseDPO [DenseDPO], extends this paradigm to the video domain by refining paired data construction and the granularity of preference labels, yielding videos with superior visual quality and motion dynamics. A primary limitation of these methods, however, is their offline nature, which precludes online parameter updates. This challenge has been recently addressed by verifiable reward-based models like OpenAI’s o1 [OpenAI-O1] and Deepseek-R1 [Deepseek-R1, shao2024deepseekmath]. Concurrently, the growing interest in online learning has spurred various online Reinforcement Learning (RL) methods, including ReFL [ReFL], RAFT [raft], DRaFT [DRaFT], FlowGRPO [flowgrpo], DanceGRPO [dancegrpo], and SRPO [SRPO], which have demonstrated promising results in enhancing image aesthetics and text-rendering fidelity. Nevertheless, their application to subject-consistent video generation remains limited. To bridge this gap, we introduce task-specific rewards tailored for subject-consistent video generation, a method that mitigates reward hacking and delivers videos with high identity consistency.
Figure 2: ID-Crafter Overview. Our model incorporates a hierarchical identity-preserving attention mechanism and a VLM that performs reasoning on the multimodal input into a video DiT to enable multi-subject video generation. An online RL stage further refines the concept alignment.
@@ 3 Method
3.1 Rectified Flow for Video Generation
Our methodology is founded upon a latent video diffusion transformer [peebles2023scalable, wan2025] trained with standard Rectified Flow (RF) [lipman2022flow, liu2022flow, esser2024scaling]. In latent video diffusion, a video clip $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is first projected into a latent space by the encoder $\mathcal{E}$ of a variational autoencoder (VAE), yielding the representation $\mathbf{z}_{0}:=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{f\times c\times h\times w}$. The corresponding VAE decoder is denoted by $\mathcal{D}(\mathbf{z}_{0})$. The generative process is formulated based on Rectified Flow, which defines a straight-line trajectory between a sample from the data distribution $\mathbf{z}_{0}\sim p_{\text{data}}$ and a sample from a simple prior distribution, typically a standard Gaussian $\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The trajectory is parameterized by a time variable $t\in[0,1]$:
$\mathbf{z}_{t}=(1-t)\mathbf{z}_{0}+t\bm{\epsilon}.$ (1)
@@ 3.1 Rectified Flow for Video Generation
$\mathcal{L}_{\text{RF}}=\mathbb{E}_{t,\mathbf{z}_{0},\bm{\epsilon}}\left[w(t)\left\lVert\mathbf{v}_{\bm{\theta}}(\mathbf{z}_{t},t,\mathbf{C}_{\text{ctx}})-(\bm{\epsilon}-\mathbf{z}_{0})\right\rVert_{2}^{2}\right],$ (2)
@@
3.2 ID-Crafter: Multi-Subject Video Generation
@@
As shown in Fig. 2, ID-Crafter extends the DiT-based latent video diffusion architecture by introducing two key architectural innovations to handle the complexities of multi-subject conditioning: (1) A hierarchical identity-preserving attention mechanism, which is designed to hierarchically aggregate features. (2) Semantic understanding via a pretrained VLM, which we employ as a sophisticated encoder for both textual and visual inputs. To further incentivize the model’s semantic alignment and identity preservation, we incorporate an online reinforcement learning phase subsequent to the initial flow matching training.
Hierarchical Attention and Semantic Grounding.
@@
Our model is based on a Wan Video [wan2025], where latent video tokens and conditioning tokens are processed together. We first encode the input reference images $\mathcal{I}$ using a pretrained image encoder to obtain visual feature maps $\{\mathbf{F}_{k}\}_{k=1}^{N}$, where each $\mathbf{F}_{k}\in\mathbb{R}^{c\times h\times w}$ corresponds to the $k$-th subject. These feature maps are then flattened into token sequences $\{\mathbf{f}_{k}\}_{k=1}^{N}$, with each $\mathbf{f}_{k}\in\mathbb{R}^{hw\times c}$. To effectively integrate information from multiple subjects and modalities, we introduce a hierarchical identity-preserving attention mechanism. This operates in three stages: (1) Intra-subject attention captures fine-grained details within each subject; (2) Inter-subject attention models the interactions between different subjects; and (3) Multi-modal attention fuses subject features with text and video tokens. This hierarchical design is crucial for preserving individual identities while capturing complex inter-subject dynamics.
Recently, VLMs have seen increasing adoption in modern video generation models [gao2025seedance, zhang2025waver], yet their potential for disentangling complex multi-subject dynamics remains significantly underutilized. Notably, no prior open-source work has integrated a VLM as the core reasoning engine within the powerful open-source Wan-Video architecture. Our approach fundamentally differs by harnessing the VLM’s granular cross-modal reasoning to actively guide our Hierarchical Identity-Preserving Attention. This transforms the VLM from a static encoder into a dynamic semantic guide. We employ the Qwen2.5-VL model [Qwen2.5-VL] to process the textual prompt $\mathbf{C}_{\text{txt}}$ and reference images $\mathcal{I}$, yielding semantically enriched tokens: $\mathbf{f}_{\text{txt}}=\text{VLM}_{\text{enc}}(\mathbf{C}_{\text{txt}},\mathcal{I})\in\mathbb{R}^{l^{\prime}\times c}$. This structured representation provides an explicit signal to the generative process. The complete set of conditioning tokens $\mathbf{C}_{\text{ctx}}$ is subsequently formed by concatenating these enhanced text tokens with the individual subject tokens: $\mathbf{C}_{\text{ctx}}=[\mathbf{f}_{\text{txt}};\mathbf{f}_{1};\ldots;\mathbf{f}_{N}]\in\mathbb{R}^{(l^{\prime}+N\cdot hw)\times c}.$ Our central innovation is thus bridging the VLM’s high-level semantic reasoning with the generative model’s low-level spatial layout control.
3.3 Post-training with Online RL
Post-training the base model with a perceptual reward for identity preservation presents a formidable challenge, primarily due to the credit assignment problem. How to adjust the intricate computations within the hierarchical ID attention layers to enhance a holistic, video-level reward? While DPO-style methods [DenseDPO, DPO] offer computational efficiency, their offline nature precludes the online parameter updates essential for this fine-tuning task. Alternatively, standard policy-gradient algorithms [ReFL, DRaFT] are often impractical in this context. They necessitate training a separate value function to estimate rewards, which is unstable and sensitive to hyperparameters. Following recent advances [flowgrpo, dancegrpo], for each condition $q$, ID-Crafter samples a group of outputs $\{o_{1},o_{2},\dots,o_{G}\}$ from the old policy $\pi_{\theta_{old}}$ and then optimizes the policy model by maximizing the following objective: